OpenStudy (anonymous):

Two identical positive charges exert a repul- sive force of 6.4×10−9N when separated by a distance 4.1×10−10m.Calculate the charge of each. The Coulomb constant is 8.98755×109N·m2/C2. Answer in units of C

OpenStudy (anonymous):

\[6.4\times10^{-9}\text{ N}=\frac{\left(8.98755\times10^{-9}\frac{\text{N}\cdot\text{m}^2}{\text{C}^2}\right)q^2}{(2.1\times10^{-10}\text{ m})^2}\] this coming from the formula for Coulomb's law, \[F=\frac{k~q_1~q_2}{d^2}\] Since the charges are identical, \(q_1=q_2\).

OpenStudy (anonymous):

i dont understand what to multiple and plug in ? this is new to me